OpenStudy (some.random.cool.kid):

Solve x2 − 5x + 1 = 0 using the quadratic formula.

OpenStudy (some.random.cool.kid):

step one is to -1 by both sides?

Nnesha (nnesha):

now just use the quadraitc formula

Nnesha (nnesha):

no**

Nnesha (nnesha):

you already know the formula so just plugin a ,b ,c values

OpenStudy (some.random.cool.kid):

yeah but first -1 to the 1 and 0 go first dont they i have to simplify them.

OpenStudy (michele_laino):

comparing your equation with the general quadratic equation: \[a{x^2} + bx + c = 0\] we get: a=1, b=-5, and c=1 now, please apply this formula: \[x = \frac{{ - b \pm \sqrt {{b^2} - 4 \times a \times c} }}{{2 \times a}}\]

Nnesha (nnesha):

no quadratic equation should be in standard form \[\huge\rm Ax^2+Bx+C=0\] equal to zero

Nnesha (nnesha):

that was for completing the square method but when we use quadratic formula we ned `standard form`

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

so what is the standard is it like ... idk actually

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha no quadratic equation should be in standard form \[\huge\rm Ax^2+Bx+C=0\] equal to zero \(\color{blue}{\text{End of Quote}}\) ^^that's the standard form

Nnesha (nnesha):

where a=leading coefficient b=middle term c=constant term

Nnesha (nnesha):

now just plugin a ,b, c values into the quadratic formula

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

writing all these down just in case your wondering... very helpful info for big exam EOC..

Nnesha (nnesha):

that's Awesome take ur time write theses down and practice as much as u can

OpenStudy (some.random.cool.kid):

k so what do i next because i wasn't sure...

Nnesha (nnesha):

did you substitute a ,b c for their values ?

OpenStudy (some.random.cool.kid):

a is x^2 right ?

OpenStudy (some.random.cool.kid):

and b is 5x?

Nnesha (nnesha):

no just the coefficient not the variable

Nnesha (nnesha):

look at the Michele'z comment a=1 bec x^2 is same as 1x^2

OpenStudy (some.random.cool.kid):

so 5 since its not a variable?

OpenStudy (some.random.cool.kid):

/so your saying 1^2? i dont get how thats possible...

Nnesha (nnesha):

yes 5 is b

Nnesha (nnesha):

there is invisible one at front of x^2 x^2 is same as 1x^2 so leading coefficient is one

Nnesha (nnesha):

i'm not saying x =1 leading coefficient is one

Nnesha (nnesha):

\[\huge\rm \color{red}{1}x^2+\color{blue}{5}x+\color{green}{1}=0\] \[\huge\rm \color{ReD}{A}x^2+\color{blue}{B}x+\color{green}{C}=0\]

OpenStudy (some.random.cool.kid):

alright

Nnesha (nnesha):

oky now just plugin

OpenStudy (some.random.cool.kid):

k thanks i gtg thanks for your help. :D